Melanie is deciding which of the two long term savings plan to choose for investing $9000: 6.2% per annum, compounding semi-annually, or 5.75% per annum compounded quarterly. If she plans to leave the money in the plan of 8 years, which is the better option?


1
Expert's answer
2021-06-16T15:04:08-0400

"A = P(1 + \frac{r}{n})^{nt}"

In the formula

  • A = Accrued amount (principal + interest)
  • P = Principal amount
  • r = Annual interest rate as a decimal
  • n = number of compounding periods per unit of time
  • t = time in decimal years

"A = P(1 + \frac{r}{n})^{nt}\\
A = 9,000.00(1 + \frac{0.062}{2})^{(2)(8)}\\
A = 9,000.00(1 + 0.031)^{(16)}\\
A = \$14,668.35"

Interest=amount-principal

"I=14668.35-9000=\$5668.35"

"A = 9,000.00(1 +\frac{ 0.0575}{4})^{(4)(8)}\\
A = 9,000.00(1 + 0.014375)^{(32)}\\
A = \$14,210.05"

Interest=amount-principal

"I=14210.05-9000=\$5210.05"

The first option is the best since its interest is high.